OpenStudy (anonymous):

Help with multiple choice anyone? The line y − 6 = 2(x + 4) passes through which one of the following points? A. (−4, −6) B. (4, −6) C. (−4, 6) D. (2, 6)

OpenStudy (pinksapphire):

First, you have to get y by itself.

OpenStudy (anonymous):

well i got c, but i need someone to check this and a few other questions, is that right?

OpenStudy (anonymous):

C is correct

OpenStudy (pinksapphire):

Yes, it is right.

OpenStudy (anonymous):

Which of the following is an equation of the line that passes through (3, −5) and has a slope of 2? A. y + 5 = 2(x − 3) B. y − 5 = 2(x + 3) C. y − 5 = 3(x − 2) D. y + 5 = 3(x − 2)

OpenStudy (anonymous):

that's another question, and i got a

OpenStudy (anonymous):

A is correct

OpenStudy (pinksapphire):

Yes, A is correct.

OpenStudy (anonymous):

thanks you! can you just check my other questions?

OpenStudy (anonymous):

A horizontal line through the point (4, 6) is defined by which one of the following equations? A. x = 4 B. x = 6 C. y = 4 D. y = 6

OpenStudy (anonymous):

i think it's d

OpenStudy (anonymous):

@PinkSapphire can you help?

OpenStudy (pinksapphire):

Sure :)

OpenStudy (pinksapphire):

Yes, D is correct.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

so i have 2 more

OpenStudy (anonymous):

What is the slope of the line given by the equation y + 2 = −3(x − 7)? A. m = −3 B. m = 3 C. m = 2 D. m = −7

OpenStudy (anonymous):

it's A

OpenStudy (pinksapphire):

A is correct.

OpenStudy (anonymous):

Which graph represents the equation y − 3.6 = 1.6(x − 1)?

OpenStudy (anonymous):

graph A

OpenStudy (anonymous):

is that right? @PinkSapphire

OpenStudy (pinksapphire):

Yes, graph A is correct.

OpenStudy (anonymous):

thank you!!!!

OpenStudy (pinksapphire):

No problem...do you need any more help?

OpenStudy (anonymous):

nope, all done! thanks again!

OpenStudy (pinksapphire):

Okay :)